SUBMILLIMETER SPECTROSCOPY OF ZnO (X^{1}Σ^{+})

\underline{L. N. ZACK}, R.L. PULLIAM, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of ZnO (X^{1}Σ^{+}) has been measured using millimeter-wave direct-absorption techniques. ZnO was produced by the reaction of zinc vapor and N_{2}O under d.c. discharge conditions. Over the frequency range 239-514 GHz, rotational transitions were measured for the ground and several vibrationally excited states of the five stable zinc isotopologues of ZnO. The ^{1}Σ^{+} ground state was confirmed, and the spectroscopic constants and equilibrium parameters were calculated. The bond length of zinc oxide was determined to be 1.708 Å.